On my HW, there as a problem where I知 stuck on it. The problem is to solve x + 2y = -5 and 4x - 3y = 2 using the elimination method and ive already simplified it to 5x - y = -3, then thats where I get stuck.

Answer by Alan3354(69443)

The problem is to solve x + 2y = -5 and 4x - 3y = 2 using the elimination method and ive already simplified it to 5x - y = -3, then thats where I get stuck.
----------------
x + 2y = -5
4x - 3y = 2
----------
To eliminate a variable, multiply the equations to make the coefficients of one of the variables equal.
For this one, I'd eliminate x.
Multiply the 1st eqn by 4:
4x + 8y = -20
4x - 3y = 2
---------------------- Subtract